OpenStudy (lilai3):

PLEASE HELP! I AM IN NEED OF HELP! PLEASE HELP! I WOULD REALLY APPRECIATE IT IF YOU WOULD HELP ME PLEASE OVER HERE...THESE STUFF IS LIKE REALLY SORTA BORING... CAN YOU PLEASE CHECK IF I AM RIGHT OR NOT???? THANK YOUS. Travel: On a map, the distance between Charleston and Columbia, South Carolina, is 5 inches. If the scale of the map is 7/8 inch= 20 miles, about how long would it take the Garcia family to drive from Charleston to Columbia if they drove 60 miles per hour?

OpenStudy (mathstudent55):

This problem has two parts. First using the map information and scale, you must find the distance between Charleston and Columbia. Then using the distance and the speed you can find the time. For the first part, use a proportion: 7/8 inch it to 20 miles as 5 inches is to d miles. \[\frac{ \frac{ 7 }{ 8 } }{ 20 } = \frac{ 5 }{ d }\] Solve the proportion with crossmultiplication: (7/8)d = 5 x 20 (7/8)d = 100 7d = 100 x 8 7d = 800 d = 800/7 = 114.3 miles Now for the second part, d = vt, or distance = velocity (speed) times time. You knoww the distance and the speed, solve for time.

OpenStudy (lilai3):

but then what do you do with the 60 miles per hour the Garcia family is going? i got a different answer...like 190 sumding...

OpenStudy (lilai3):

and you said this: Solve the proportion with crossmultiplication: (7/8)d = 5 x 20 (7/8)d = 100 7d = 100 x 8 7d = 800 d = 800/7 = 114.3 miles but where did the "times 8" come from?

OpenStudy (mathstudent55):

d = vt 114.3 = 60t Solve for t by dividing both sides by 60 114.3/60 = 60t/60 t = 1.90 hours Since 1.90 hours is not c a very common way of expressing a time, change in into 1 hour + 0.9 hour Since there are 60 minutes in an hour, 1 hour + 0.9 hour ( (60 minutes) / (1 hour) ) You get 1 hour and 54 minutes

OpenStudy (lilai3):

ohh i get it now... it makes more sense. thank you so much, but @mathstudent55, will you still be able to help me for one more problem please?

OpenStudy (mathstudent55):

The 8 came from here: \[\frac{ 7 }{ 8 } d = 100\] Multiply both sides by 8 \[\frac{ 7 }{ 8 } d \times 8 = 100 \times 8\] On the left side, the 8's cancel out

OpenStudy (mathstudent55):

yes, next problem

OpenStudy (lilai3):

oh. okay :D thanks i gave you a medal for helping me. you're awesome!

OpenStudy (mathstudent55):

Post the other problem as a new problem and I'll try to help you

OpenStudy (lilai3):

Animals: An average adult giraffe is 18 feet tall. A newborn giraffe is about 6 feet tall. Kayla is building a model of a mother giraffe and her newborn. She wants the model to be no more than 17 inches high. Choose an appropriate scale for a model of the giraffes. Then use it to find the height of the mother and the height of the newborn giraffe. I got to admit, this one is a little bit harder...LOLs. :D Thanks for helping me.

OpenStudy (mathstudent55):

You need to turn a real length of 18 ft into a model length of at most 17 inches. First let's look at that. Since 1 ft = 12 in., 18 ft = 216 in.

OpenStudy (mathstudent55):

A scale of 12:1 would seem to be good, you'd make the 18 ft giraffe into a 18 inch model, but you can't have a model more than 17 inches long.

OpenStudy (lilai3):

okay..so that would mean the the model of the giraffe should be up to 16 inches, but that cannot be divisible by 3... so maybe 15 inches for the adult giraffe and 5 inches for the baby giraffe? Does that work?

OpenStudy (mathstudent55):

Actually that scale would be called 1:12 (1 model size to 12 real size), but it still makes a model 18 inches long. You need a scale that will reduce the size of the model more.

OpenStudy (lilai3):

ion't getit

OpenStudy (mathstudent55):

Your suggestion is not bad, but that would mean a scale of 1:14.4 Maybe an adult giraffe could be 12 inches and the baby 4 inches

OpenStudy (lilai3):

that would be 1:3 right?

OpenStudy (mathstudent55):

Since the adult giraffe measures 18 ft = 216 inches, a 12 inch model would be 216/12 = 18 times smaller. Then a scale of 1:18 could work

OpenStudy (mathstudent55):

It's not a scale of 1:3. Remember the 1 : 3 here is the difference between the newborn (6 ft) and the adult (18 ft)

OpenStudy (lilai3):

oh right. but Kayla wants it less than 17 inches. i know 18 would work for certain but like if it was 19 it coulda been bettah

OpenStudy (mathstudent55):

The largest model you can make is 17 inches long. You need to use a length for the mother that is (1) less than 17 inches and (2) also divisible by three and (3) also that you can write its scale easily.

OpenStudy (mathstudent55):

Your idea of using 15 inches for the mother and 5 for the baby are good for (1) and (2) above, but not for (3). If you use 12 inches for the mother, then you have (1) under 17 inches, (2) 12 IS divisible by 3, and (3) the scale is easily stated as 1 inch = 1.5 ft

OpenStudy (lilai3):

i saw ur comment up above saying to start a new post, and okay sure, i'll tag you in it. just wait a sec. i havta cloz dis quesshun

OpenStudy (mathstudent55):

ok

OpenStudy (lilai3):

hi. i am really, really sorry i havta leave you cuz i really need help with this, and my friggin parents are telling me to go to sleep and rite now it's like 11:22 PM and it's like really, really late and this is the first time my parents noticed i really need help with this and they don't let me stay up ...omg this is like the first time i stayed up dis l8! my friends are always sleeping at like one. I want to too, i'm not sleepy, but you know how parents are...dammit. (srry for the cursing. i had to let it out.) anyways, i really really appreciate your help and so i would like to ask you if you could just please type everything: the answer, explain how you got it, in proportions, and such. (etc.) listen, i really apppreciate it but i gotta go to school tomorrow and so can you just type everythign down here and tomorrow morning, i'll just check it and reply at around 7:30 PM -ish at Califoria USA. i got to go. bye thank you so much!

OpenStudy (lilai3):

and right now it's 11:26 PM 11/26/12. :D

OpenStudy (mathstudent55):

Read the end of the previous post. I finished the problem there. The scale is 1 inch = 1.5 feet, and the mother giraffe is 12 inches and the newborn 4 inches. Good night